OpenStudy (anonymous):

whati s “400 pounds per 340 square feet” as a rate in simplest form

OpenStudy (alexandervonhumboldt2):

\[400 lb/340ft^2\] first find the GCF. it is 20. So now divide both 400 and 340 by 20. 400/20=? 340/20=?

OpenStudy (anonymous):

400/20=20 340/20=17 so would it be 20 /17 or 17/20

OpenStudy (alexandervonhumboldt2):

it will be 20/17 as the initional fraction is 400/340

OpenStudy (anonymous):

thanks

OpenStudy (alexandervonhumboldt2):

nope)